# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v3

@@ C.3. Upper bound for $\text{IVM}^{+}[Q]$
The case of $\text{IVM}^{+}[Q]$: Let $(\mathcal{T},\chi)$ be a tree decomposition of $Q$ whose fractional hypertree width is ${\mathsf{w}(Q)}$. (Recall notation from Section B.) Fix an orientation of $\mathcal{T}$ starting from some root node $r\in V(\mathcal{T})$. Given a node $t\in V(\mathcal{T})$, we will use $\text{\sf parent}(t),\text{\sf children}(t),\text{\sf ancestors}(t)$ and $\text{\sf decendants}(t)$ to refer to the parent, children, ancestors and descendants of $t$ in $\mathcal{T}$, respectively, under the fixed orientation of $\mathcal{T}$. For every node $t\in V(\mathcal{T})$, let $\bm{X}_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}\chi(t)$ be the variables of $t$ and $Q_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}Q_{\bm{X}_{t}}$ be the restriction of $Q$ to $\bm{X}_{t}$, given by Definition B.3. We call $Q_{t}$ a bag query. Let $\bm{X}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf vars}(Q)$. By definition, the query $Q$ is equivalent to the join of the outputs of all bag queries:
@@ C.4. Upper bound for $\text{IVM}^{+}_{\delta}[Q]$
Now suppose we have an insert of a tuple $\bm{t}$ into a relation $R(\bm{Z})$ and we want to enumerate the new output tuples $\delta Q$ that are added due to this insert. We find the node $t$ in $\mathcal{T}$ where $\bm{Z}\subseteq\chi(t)$. The existence of such a node $t$ is guaranteed by the definition of a tree decomposition (Definition B.2). Now, we re-orient the tree $\mathcal{T}$ so that $t$ becomes the root. We enumerate output tuples $\bm{x}_{t}$ of $Q^{\prime\prime}_{t}$ that join with $\bm{t}$. For each such tuple $\bm{x}_{t}$, we enumerate the corresponding tuples $\bm{x}_{c}$ from $Q^{\prime\prime}_{c}$ for every child $c$ of $t$. (These are the children of $t$ under the new orientation of the tree $\mathcal{T}$). By (34), we guarantee that for each tuple $\bm{x}_{t}$ and each child $c$, there must be at least one tuple $x_{c}$ in $Q^{\prime\prime}_{c}$ that joins with $\bm{x}_{t}$. We continue this process inductively all the way to the leaves of $\mathcal{T}$. ∎
Example C.3 (for Theorem 3.1).
Consider again the query (9) from Example 5.3. Suppose now that we want to extend our solution from Example 5.3 to the $\text{IVM}^{+}_{\delta}[Q]$ problem. In addition to $Q_{1},P_{1}$ and $Q_{2}^{\prime}$ that are defined in Example 5.3, the proof of Theorem 3.1 for $\text{IVM}^{+}_{\delta}[Q]$ says that we also need to maintain the following relations:
$P_{1}^{\prime\prime}(B,C)$ $=Q_{2}^{\prime}(B,C,D)$
@@ C.5. Lower bound for $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$
In this section, we prove Theorem 3.3, which gives a lower bound on the amortized updated time for both $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ in terms of a lower bound on the static evaluation of $Q$.
Theorem 3.3.
For any query $Q$, neither $\text{IVM}^{+}[Q]$ nor $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{\omega}(Q)-1-\gamma})$ amortized update time, for any constant $\gamma>0$.
Proof of Theorem 3.3.
Fix a query $Q$. We first consider $\text{IVM}^{+}[Q]$. Suppose, for sake of contradiction, that there is a constant $\gamma>0$ where $\text{IVM}^{+}[Q]$ can be solved with amortized update time $\widetilde{\mathcal{O}}(N^{\alpha})$ where $\alpha\stackrel{{\scriptstyle\text{def}}}{{=}}\omega(Q)-1-\gamma$. We will show how to use an oracle for $\text{IVM}^{+}[Q]$ to solve $\text{Eval}[Q]$ in time $\widetilde{\mathcal{O}}(|\mathcal{D}|^{\omega(Q)-\gamma}+|Q(\mathcal{D})|)$ on any database instance $\mathcal{D}$. This contradicts the definition of $\omega(Q)$ (Definition 2.2).
We construct a stream of inserts of length $N=|\mathcal{D}|$ where we insert one tuple of $\mathcal{D}$ at a time. We apply these inserts one by one in a total time of $\widetilde{\mathcal{O}}(N^{1+\alpha})$. After the last insert, we invoke constant-delay enumeration of $Q(\mathcal{D})$. Thus, the overall time is $\widetilde{\mathcal{O}}(N^{1+\alpha}+|Q(\mathcal{D})|)$.
@@ Appendix D Missing Details in Section 6
In this section, we review some necessary background on intersection joins in the static setting. Let $Q$ be a query, $\overline{Q}$ be its time extension, and $\widehat{Q}$ be its multivariate extension. Suppose we want to evaluate $\overline{Q}$ over a given database instance $\overline{\mathcal{D}}$. This evaluation problem can be reduced (Khamis et al., 2022) to the evaluation of $\widehat{Q}$ over some database instance, $\widehat{\mathcal{D}}$, that satisfies $|\widehat{\mathcal{D}}|=\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$ and can be constructed in time $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$. Moreover, this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown (Khamis et al., 2022) by a backward reduction from each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail below. Recall that both reductions are in the context of static evaluation. However later on in Section E, we will rely on both reductions in order to prove upper and lower bounds on the IVM problem in the insert-delete setting.
D.1. A reduction from $\overline{Q}$ to $\widehat{Q}$
The reduction in (Khamis et al., 2022) from a time extension query $\overline{Q}$ to the corresponding multivariate extension query $\widehat{Q}$ relies on constructing a segment tree for the interval variable $[Z]$ in $\overline{Q}$.
@@ D.1. A reduction from $\overline{Q}$ to $\widehat{Q}$
Figure 10 depicts a segment tree $\mathcal{T}_{8}$ for $N=8$. Following (Khamis et al., 2022), we identify each node of a segment tree by a bitstring: The root is identified by the empty string $\varepsilon$; for each inner node that is identified by a bitstring $b$, the left and right child nodes are identified by the bitstrings $b\circ 0$ and respectively $b\circ 1$ (where $\circ$ denotes string concatenation). We denote by $V(\mathcal{T}_{N})$ the set of nodes of the segment tree $\mathcal{T}_{N}$ and by $\text{\sf parent}(v)$ the parent node of a node $v$. Given an interval $x\subseteq[N]$, its canonical partition with respect to $N$ is defined as the set of maximal nodes $y$ of $\mathcal{T}_{N}$ whose corresponding intervals are contained in $x$, i.e.
@@
For any interval $x\subseteq[N]$, $\text{\sf CP}_{N}(x)$ is partition of $x$; see (Khamis et al., 2022). Moreover $|\text{\sf CP}_{N}(x)|=\mathcal{O}(\log N)$. For example, in the segment tree $\mathcal{T}_{8}$ from Figure 10, we have,
@@
[…]
A segment tree $\mathcal{T}_{8}$. Each node is identified by a bitstring.
Lemma D.1 ((Khamis et al., 2022)).
@@
Theorem D.5 (Implicit in (Khamis et al., 2022)).
@@
D.2. A reduction from $\widehat{Q}$ to $\overline{Q}$
In the previous Section D.1, we saw how to reduce the static evaluation of $\overline{Q}$ to $\widehat{Q}$. We now summarize the backward reduction from the static evaluation of each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ back to $\overline{Q}$, based on (Khamis et al., 2022). Later on, we will use this backward reduction to prove lower bounds on the IVM problem in Section E.4.
Example D.6.
Continuing with Example 6.3, let’s take the query $\widehat{Q}_{123}$ from (37). Let $\widehat{\mathcal{D}}_{123}^{\prime}$ be an arbitrary database instance for $\widehat{Q}_{123}$. (Unlike Example 6.3, here we cannot make any assumption about how $\widehat{\mathcal{D}}_{123}^{\prime}$ was constructed.) We show how to use an oracle for the intersection query $\overline{Q}_{\triangle}$ from (11) in order to compute $\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123}^{\prime})$ in the same time. WLOG we can assume that each value $x$ that appears in $\widehat{\mathcal{D}}_{123}^{\prime}$ is a bitstring of length exactly $\ell$ for some constant $\ell$. (If a bitstring has length less than $\ell$, we can pad it with zeros.) We construct a segment tree $\mathcal{T}_{N^{\prime}}$ where $N^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}2^{3\ell}$, i.e. a segment tree of depth $3\ell$. (We chose $3$ in this example because it is the number of atoms in $\widehat{Q}_{123}$.) Each node $v$ in the segment tree corresponds to an interval that is contained in $[N^{\prime}]$ and is identified by a bitstring of length at most $3\ell$. Given a bitstring $b$ of length at most $3\ell$, let $\text{\sf seg}_{N^{\prime}}(b)$ denote the corresponding interval in the segment tree $\mathcal{T}_{N^{\prime}}$. We construct the following database instance $\overline{\mathcal{D}}^{\prime}$ for $\overline{Q}_{\triangle}$:
Following (Khamis et al., 2022), we can show that there is a one-to-one mapping between the output tuples of $\overline{Q}_{\triangle}(\overline{\mathcal{D}}^{\prime})$ and the output tuples of $\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123}^{\prime})$. This follows from the observation that the three intervals $\text{\sf seg}_{N^{\prime}}(z_{1}^{R})$, $\text{\sf seg}_{N^{\prime}}(z_{1}^{S}\circ z_{2}^{S})$, $\text{\sf seg}_{N^{\prime}}(z_{1}^{T}\circ z_{2}^{T}\circ z_{3}^{T})$ overlap if and only if $z_{1}^{R}=z_{1}^{S}=z_{1}^{T}$ and $z_{2}^{S}=z_{2}^{T}$. (Recall that the bitstrings $z_{1}^{R},z_{1}^{S},\ldots$ all have the same length $\ell$.) Moreover note that $\overline{R}^{\prime},\overline{S}^{\prime},\overline{T}^{\prime}$ have the same sizes as $\widehat{R}_{123}^{\prime},\widehat{S}_{123}^{\prime},\widehat{\mathcal{T}}_{123}^{\prime}$ respectively and they can be constructed in linear time.
Extrapolating from the above example, we introduce the following concept:
@@ D.2. A reduction from $\widehat{Q}$ to $\overline{Q}$
Theorem D.9 (Implicit in (Khamis et al., 2022)).
@@ Appendix E Missing Details from Section 7
Theorem 3.4 gives an upper bound on the amortized update time for both the $\text{IVM}^{\pm}$ and $\text{IVM}^{\pm}_{\delta}$ problems in terms of $N$, the length of the update stream. Corollary 3.5 tightens this upper bound from $N$ to the current size of the database, $|\mathcal{D}^{(\tau)}|$. In contrast, Theorem 3.6 gives a lower bound on the amortized update time only for $\text{IVM}^{\pm}_{\delta}$ (in terms of $N$ and by extension $|\mathcal{D}^{(\tau)}|$). We prove Theorem 3.4 for $\text{IVM}^{\pm}$ in Section E.1 and for $\text{IVM}^{\pm}_{\delta}$ in Section E.2. We prove Corollary 3.5 in Section E.3. In Section E.4, we prove Theorem 3.6 (for $\text{IVM}^{\pm}_{\delta}$).
@@ E.1. Upper Bound for $\text{IVM}^{\pm}[Q]$
Example 7.1 introduces the high-level idea of our algorithm for $\text{IVM}^{\pm}[Q]$. Algorithm 2 summarizes our algorithmic framework for $\text{IVM}^{\pm}[Q]$ that meets Theorem 3.4. Algorithm 2 assumes that the length $N$ of the update stream is known in advance. However, the following remark explains how to remove this assumption.
@@
For the rest of this section, we prove that Algorithm 2 meets Theorem 3.4. Fix a query $Q$ together with its time extension $\overline{Q}$ (Definition 6.2) and its multivariate extension $\widehat{Q}$ (Eq. (3)). We split the proof of Theorem 3.4 for $\text{IVM}^{\pm}[Q]$ into three parts. First, we reduce the maintenance of $Q$ to the maintenance of $\overline{Q}$ (Lemma E.1). Then, we reduce the latter problem to the maintenance of $\widehat{Q}$ (Lemma E.3). Finally, we show that $\widehat{Q}$ can be maintained with $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}-1})$ amortized update time, where ${\mathsf{w}(\widehat{Q})}$ is the fractional hypertree width of $\widehat{Q}$ (Lemma E.4). Theorem 3.4 for $\text{IVM}^{\pm}[Q]$ then follows from Lemmas E.1, E.3, and E.4.

### v4

@@ C.3. Upper bound for $\text{IVM}^{+}[Q]$
The case of $\text{IVM}^{+}[Q]$: Let $(\mathcal{T},\chi)$ be a tree decomposition of $Q$ whose fractional hypertree width is ${\mathsf{w}(Q)}$. (Recall notation from Section B.) Fix an orientation of $\mathcal{T}$ starting from some root node $r\in V(\mathcal{T})$. Given a node $t\in V(\mathcal{T})$, we will use $\text{\sf parent}(t),\text{\sf children}(t),\text{\sf ancestors}(t)$ and $\text{\sf decendants}(t)$ to refer to the parent, children, ancestors and descendants of $t$ in $\mathcal{T}$, respectively, under the fixed orientation of $\mathcal{T}$. For every node $t\in V(\mathcal{T})$, let $\bm{X}_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}\chi(t)$ be the variables of $t$ and $Q_{t}\stackrel{{\scriptstyle\text{def}}}{{=}}Q_{\bm{X}_{t}}$ be the restriction of $Q$ to $\bm{X}_{t}$, given by Definition 2.2. We call $Q_{t}$ a bag query. Let $\bm{X}\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf vars}(Q)$. By definition, the query $Q$ is equivalent to the join of the outputs of all bag queries:
@@ C.4. Upper bound for $\text{IVM}^{+}_{\delta}[Q]$
Now suppose we have an insert of a tuple $\bm{t}$ into a relation $R(\bm{Z})$ and we want to enumerate the new output tuples $\delta Q$ that are added due to this insert. We find the node $t$ in $\mathcal{T}$ where $\bm{Z}\subseteq\chi(t)$. The existence of such a node $t$ is guaranteed by the definition of a tree decomposition (Definition 2.1). Now, we re-orient the tree $\mathcal{T}$ so that $t$ becomes the root. We enumerate output tuples $\bm{x}_{t}$ of $Q^{\prime\prime}_{t}$ that join with $\bm{t}$. For each such tuple $\bm{x}_{t}$, we enumerate the corresponding tuples $\bm{x}_{c}$ from $Q^{\prime\prime}_{c}$ for every child $c$ of $t$. (These are the children of $t$ under the new orientation of the tree $\mathcal{T}$). By (35), we guarantee that for each tuple $\bm{x}_{t}$ and each child $c$, there must be at least one tuple $x_{c}$ in $Q^{\prime\prime}_{c}$ that joins with $\bm{x}_{t}$. We continue this process inductively all the way to the leaves of $\mathcal{T}$. ∎
@@ C.5. Lower bound for $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$
In this section, we prove Proposition 4.3, which gives a lower bound on the amortized updated time for both $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ in terms of a lower bound on the static evaluation of $Q$.
Proposition 4.3.
For any query $Q$ and any constant $\gamma>0$, neither $\text{IVM}^{+}[Q]$ nor $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{\omega}(Q)-1-\gamma})$ (amortized) update time.
Proof of Proposition 4.3.
Fix a query $Q$. We first consider $\text{IVM}^{+}[Q]$. Suppose, for sake of contradiction, that there is a constant $\gamma>0$ where $\text{IVM}^{+}[Q]$ can be solved with amortized update time $\widetilde{\mathcal{O}}(N^{\kappa})$ where $\kappa\stackrel{{\scriptstyle\text{def}}}{{=}}\omega(Q)-1-\gamma$. We will show how to use an oracle for $\text{IVM}^{+}[Q]$ to solve $\text{Eval}[Q]$ in time $\widetilde{\mathcal{O}}(|\mathcal{D}|^{\omega(Q)-\gamma}+|Q(\mathcal{D})|)$ on any database instance $\mathcal{D}$. This contradicts the definition of $\omega(Q)$ (Definition 3.2).
We construct a stream of inserts of length $N=|\mathcal{D}|$ where we insert one tuple of $\mathcal{D}$ at a time. We apply these inserts one by one in a total time of $\widetilde{\mathcal{O}}(N^{1+\kappa})$. After the last insert, we invoke constant-delay enumeration of $Q(\mathcal{D})$. Thus, the overall time is $\widetilde{\mathcal{O}}(N^{1+\kappa}+|Q(\mathcal{D})|)$.
@@ Appendix D Missing Details in Section 6
In this section, we review some necessary background on intersection joins in the static setting. Let $Q$ be a query, $\overline{Q}$ be its time extension, and $\widehat{Q}$ be its multivariate extension. Suppose we want to evaluate $\overline{Q}$ over a given database instance $\overline{\mathcal{D}}$. This evaluation problem can be reduced (Abo Khamis et al., 2022) to the evaluation of $\widehat{Q}$ over some database instance, $\widehat{\mathcal{D}}$, that satisfies $|\widehat{\mathcal{D}}|=\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$ and can be constructed in time $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|)$. Moreover, this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown (Abo Khamis et al., 2022) by a backward reduction from each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail below. Recall that both reductions are in the context of static evaluation. However later on in Section E, we will rely on both reductions in order to prove upper and lower bounds on the IVM problem in the insert-delete setting.
D.1. Reduction from $\overline{Q}$ to $\widehat{Q}$
The reduction in (Abo Khamis et al., 2022) from a time extension query $\overline{Q}$ to the corresponding multivariate extension query $\widehat{Q}$ relies on constructing a segment tree for the interval variable $[Z]$ in $\overline{Q}$.
@@ D.1. Reduction from $\overline{Q}$ to $\widehat{Q}$
Figure 2 depicts a segment tree $\mathcal{T}_{8}$ for $N=8$. Following (Abo Khamis et al., 2022), we identify each node of a segment tree by a bitstring: The root is identified by the empty string $\varepsilon$; for each inner node that is identified by a bitstring $b$, the left and right child nodes are identified by the bitstrings $b\circ 0$ and respectively $b\circ 1$ (where $\circ$ denotes string concatenation). We denote by $V(\mathcal{T}_{N})$ the set of nodes of the segment tree $\mathcal{T}_{N}$ and by $\text{\sf parent}(v)$ the parent node of a node $v$. Given an interval $x\subseteq[N]$, its canonical partition with respect to $N$ is defined as the set of maximal nodes $y$ of $\mathcal{T}_{N}$ whose corresponding intervals are contained in $x$, i.e.
@@
For any interval $x\subseteq[N]$, $\text{\sf CP}_{N}(x)$ is partition of $x$; see (Abo Khamis et al., 2022). Moreover $|\text{\sf CP}_{N}(x)|=\mathcal{O}(\log N)$. For example, in the segment tree $\mathcal{T}_{8}$ from Figure 2, we have,
@@
Lemma D.1 ((Abo Khamis et al., 2022)).
@@
Theorem D.5 (Implicit in (Abo Khamis et al., 2022)).
@@
D.2. Reduction from $\widehat{Q}$ to $\overline{Q}$
In the previous Section D.1, we saw how to reduce the static evaluation of $\overline{Q}$ to $\widehat{Q}$. We now summarize the backward reduction from the static evaluation of each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ back to $\overline{Q}$, based on (Abo Khamis et al., 2022). Later on, we will use this backward reduction to prove lower bounds on the IVM problem in Section E.4.
Example 6.5 introduces the high-level idea of the backward reduction. Extrapolating from Example 6.5, we introduce the following concept:
@@ D.2. Reduction from $\widehat{Q}$ to $\overline{Q}$
Theorem D.8 (Implicit in (Abo Khamis et al., 2022)).
@@ Appendix E Missing Details from Section 7
Theorem 4.4 gives an upper bound on the amortized update time for both the $\text{IVM}^{\pm}$ and $\text{IVM}^{\pm}_{\delta}$ problems in terms of the current database size $|\mathcal{D}|$. In order to prove it, we first prove Lemma 7.1 which gives a weaker upper bound where $|\mathcal{D}|$ is replaced by the number of single-tuple updates $N$. We prove Lemma 7.1 for $\text{IVM}^{\pm}$ in Section E.1 and for $\text{IVM}^{\pm}_{\delta}$ in Section E.2. In Section E.3, we show how to use Lemma 7.1 as a black box in order to prove Theorem 4.4.
In contrast to Theorem 4.4, Theorem 4.5 gives a lower bound on the amortized update time only for $\text{IVM}^{\pm}_{\delta}$ in terms of the current database size $|\mathcal{D}|$. We prove it in Section E.4, by proving a stronger lower bound where $|\mathcal{D}|$ is replaced by the number of single-tuple updates $N$.
@@ E.1. Upper Bound for $\text{IVM}^{\pm}[Q]$
Example 7.2 introduces the high-level idea of our algorithm for $\text{IVM}^{\pm}[Q]$. Algorithm 2 summarizes our algorithmic framework for $\text{IVM}^{\pm}[Q]$ that meets Lemma 7.1. Algorithm 2 assumes that the length $N$ of the update stream is known in advance. However, the following remark explains how to remove this assumption.
@@
For the rest of this section, we prove that Algorithm 2 meets Lemma 7.1. Fix a query $Q$ together with its time extension $\overline{Q}$ (Definition 6.2) and its multivariate extension $\widehat{Q}$ (Eq. (7)). We split the proof of Lemma 7.1 for $\text{IVM}^{\pm}[Q]$ into three parts. First, we reduce the maintenance of $Q$ to the maintenance of $\overline{Q}$ (Lemma E.1). Then, we reduce the latter problem to the maintenance of $\widehat{Q}$ (Lemma E.3). Finally, we show that $\widehat{Q}$ can be maintained with $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}-1})$ amortized update time, where ${\mathsf{w}(\widehat{Q})}$ is the fractional hypertree width of $\widehat{Q}$ (Lemma E.4). Lemma 7.1 for $\text{IVM}^{\pm}[Q]$ then follows from Lemmas E.1, E.3, and E.4.